# Global parameterizations of $\pi\pi$ scattering with dispersive constraints: beyond the S0 wave

### v1

@@ Abstract
We provide new global parameterizations of $\pi\pi\to\pi\pi$ scattering for the S2, P, D, F, and G partial waves up to at least 1.8 GeV, easy to implement for phenomenological use. With earlier S0-wave parameterizations, slightly updated here, they reproduce previous partial wave dispersion analyses up to the $\pi\omega$ threshold. In addition, these new parameterizations have improved their description of recent P-wave data, the inelasticity in various waves, and their fulfillment of Roy-like and forward dispersion relations. The latter now test very high partial waves and have an improved matching with the Regge regime, extending those with P-wave contributions up to 1.6 GeV. Above 1.6 GeV and up to 1.8 GeV, or sometimes somewhat beyond, the parameterizations are simple unconstrained fits to data.
@@ I Introduction
Still, modern Hadron Physics demands precise and model-independent meson-meson scattering parameterizations. This requires the use of dispersion relations, which over roughly the last two decades have been successfully applied to describe scattering data on $\pi\pi$ [50, 51, 52, 53, 54, 55, 56, 57, 58, 59], but also on $\pi N$ [60, 61], $\gamma\gamma\to\pi\pi$ [62, 63], $\pi K$ [64, 65] and $\pi\pi\to K\bar{K}$ [66] (see [67] for a review of the last two).
@@
Unfortunately, such rigorous dispersive results, are not always suitable for later practical use. There are several reasons for this. First, they are often obtained numerically from integral equations, which makes them inconvenient for further phenomenological or experimental analyses. Second, they are frequently parameterized in the real axis by piecewise functions. They provide flexibility, but cannot be directly continued to the complex plane in search of poles. Finally, of the usual constraints used in the literature, partial-wave dispersion relations are in practice limited to energies around 111\,1GeV for $\pi\pi$ and $\pi K$ scattering, and to 1.41.41.4\,1.4 GeV for forward dispersion relations. Above these energies, partial-wave experimental data have not been tested nor described with dispersive constraints. In this work, we will partially alleviate these caveats.
@@
Recently, two of us with another collaborator [68] provided a set of relatively simple and ready-to-use “global” parameterizations that describe the scalar-isoscalar (S0) and vector-isovector (P1 or just P) partial waves of $\pi\pi\to\pi\pi$ scattering data up to somewhat below 2 GeV. They mimic the central values and uncertainties of the dispersive data analysis in [54], which was dispersively constrained up to 1.42 GeV. Beyond this energy, these parameterizations were three purely phenomenological fits to three different data sets from the literature. In addition, being analytic expressions, these parameterizations can be evaluated in the complex plane and they were made to reproduce the dispersive results, including the poles associated with the $\sigma/f_{0}(500)$, $f_{0}(980)$ and $\rho(770)$ resonances. Later, it was found in [69] that the poles $f_{0}(1370)$ and $f_{0}(1500)$ are also present in the dispersive analysis and global parameterization of the S0 wave. Furthermore, these global parameterizations are also consistent with the threshold parameters and the value of the S0-wave Adler zero in [54]. They have been widely used by both the theoretical and experimental hadron and particle physics communities.
@@
The aim of this work is threefold. First, extend this “global” parameterization approach beyond the partial waves S0 and P up to those with angular momentum $\ell=4$, trying to avoid piece-wise parametrizations whenever possible. Second, improve the treatment of the P-wave inelasticity, allowing it to start at the $\pi\omega$ threshold. This was neglected in [54, 68] [^1] We thank G. Colangelo, M. Hoferichter, and P. Stoffer for pointing this to us., which focused more on the S0 wave and its resonances, but there is a considerable interest in the precise description of this wave. Third, extend the forward dispersion relation constraints to higher energies, while improving the matching with the high-energy Regge description. These improvements will have a relatively small effect on the S0 wave, whose global parameters will be slightly updated, but only because of the indirect effects of the changes in the other waves.
@@
The plan of this work is as follows: Sect. II presents the data to be described. In Sect. III we present the partial-wave parameterizations. In Sect. IV we will revisit the dispersive constraints and obtain constrained Global Fits. We will discuss and summarize our results in Sect. V.
@@ II The input to be described
There are several data sets on $\pi\pi\to\pi\pi$ scattering partial waves $t^{I}_{\ell}(s)$, which almost reach a center-of-mass energy of $\sqrt{s}\sim 2\,$ GeV [20, 21, 22, 23, 24, 25, 26, 27, 28]. Here $I=0,1,2$ stands for isospin and $\ell=0,1,2,3,4...$ for angular momentum, although we will follow the usual spectroscopic notation and refer to them as S, P, D, F, G… waves, followed by their isospin number.
Unfortunately, these are indirect measurements, extracted as a sub-process of $\pi N\to\pi\pi N^{\prime}$ scattering, which is the observed reaction. This analysis assumes the dominance of an almost on-shell one-pion exchange, along with other approximations. As a consequence of this indirect extraction, $\pi\pi$ data sets have large systematic uncertainties, and are often incompatible with each other, even when extracted from the same $\pi N\to\pi\pi N^{\prime}$ experiment. Furthermore, simple fits to individual data sets or to averaged sets do not comply well with dispersion relations [70, 71, 72, 73, 54]. This is why such relations were used to eliminate inconsistent data sets or as constraints to obtain a Constrained Fit to Data (CFD) [54]. In addition, these CFDs fulfill the normality requirements of the residual distribution [74], thus allowing for standard error-propagation methods. Consequently, the CFD parameterizations will be part of our input.
Why, then, not to use the CFD directly? Actually, it is perfectly fine to use them, although they were constructed with piecewise functions and only up to energies around 1.41.41.4\,1.4GeV. Moreover, the focus of [54] was the low-energy parameters and the region of the $\sigma/f_{0}(500)$, $f_{0}(980)$, and $\rho(770)$ resonances. The description above 1.1 GeV, or of higher waves, was provided primarily to serve as input for the integral representation of the S0 and P amplitudes below that energy. The need for “global” parameterizations is justified in that many applications require a wider energy range and waves with higher angular momentum, and because the CFD piecewise expressions cannot be straightforwardly continued to the complex plane to describe resonance poles. This is why in [68] some of us provided such global parameterizations for the S0 and P waves. Here we extend this approach to six more waves: S2, D0, D2, F, G0, and G2. In addition, we improve some details of the P wave above 0.9 GeV, expanding the dispersive constraints further than 1.4 GeV. Moreover, above roughly 1 GeV, we improve the precision of all waves beyond the S0.
@@
Hence, we will use as input the threshold parameters obtained from the CFD parameterizations in [54] or sum rules [73]. In addition, we will consider the CFD in the real $s$ axis as input for the S2 and D waves, although only below 0.9 GeV, as we plan to improve the treatment of their inelasticities, which in [54] were only taken into account above the $K\bar{K}$ threshold.
@@
Concerning the S0 wave for energies above 1.4 GeV, “Solution b” and “Solution (- - -)” are very incompatible. Other solutions were already disfavored in the very same CERN-Munich 1975 analysis, although the “(- + -) solution” has recently been resurrected [78]. Nevertheless, all these solutions have caveats (see [79, 45] and [70] for detailed discussions). In particular, at high energies, the elastic cross section of “Solution b” is larger than the inelastic one, which is at odds with the observations in $\pi N$, $KN$, and $NN$ scattering, and there is no obvious reason why $\pi\pi$ scattering should be different. In addition, right above 1 GeV, the S0-wave data of the other CERN-Munich solutions prefer less inelasticity, i.e. the so-called “non-dip” scenario [54], which is disfavored by Roy-like dispersive constraints [54, 56]. Later on, we will remove such data points for these solutions. Moreover, if the inelasticity is large, the solution in terms of phase shift and elasticity is not unique [80, 81]. “Solution b” is an example of an almost elastic case and “Solution (- - -)” is strongly inelastic. This, of course, applies to other waves. Furthermore, a slight modification of the “(- + -) solution” by one of the members of the CERN-Munich collaboration [78] considering an inelastic S2 wave, seems consistent with the dispersive representation [68], finding some qualitative agreement with the GAMS experiment on $\pi^{-}p\rightarrow\pi^{0}\pi^{0}n$ [82]. Such a “(- + -) solution” differs from the “b” and “(- - -)” solutions only above 1.4 GeV and hints at the presence of the $f_{0}(1500)$ resonance, not evident in other solutions. Last but not least, the convergence of the partial-wave expansion is questionable, since already at 1.7 GeV the F wave is as large as the P wave, the D0 as the S0, and the D2 is larger than the S2.
@@
Not only the S0 but also the P, F, and D0 waves, have three solutions, although they are not so different among themselves. The existence of different data sets leads to three different constrained global fits in [68], almost identical up to 1.43 GeV. Namely, above 1.43 GeV: i) “Solution I” describes the data of [20, 25, 28], ii) “Solution II” describes the (- - -) data of [26], and iii) “Solution III” describes the data of the “(- + -) solution” updated in [78]. We will follow a similar strategy here, considering the three sets of data called Solutions I, II, and III again, and their corresponding fits that we will call Global Fits I, II, and III. There are no scattering data for the G0 wave and we will build an educated guess from other information.
@@
Concerning the $I=2$ waves, we also use CFD input below 0.9 GeV for the repulsive S2, and D2 often called “exotic waves”. No CFD was available for the G2 wave, which is also exotic. Above that energy we consider the 1973 data from the Rochester collaboration obtained at Brookhaven (Cohen et al. [23]) and the Paris-Bari collaboration obtained in 1973 at CERN (Durusoy et al. [21]), the 1974 CERN-Saclay data (Losty et al. [22]) and the 1977 Amsterdam-CERN-Munich data (Hoogland et al. [27]). These experiments are roughly in agreement, partly because these waves are small and the uncertainties become comparatively large. Thus, we do not contemplate alternative data sets for these waves and there will be only one unconstrained fit for each $I=2$ wave. However, since the dispersive constraints affect all waves simultaneously, the alternative data sets from other waves will indirectly affect the $I=2$ waves, which will also have three different constrained fits, although they will look rather similar in practice.
@@
Having fits for all these waves up to at least 1.8 GeV, and sometimes even up to 2.1 GeV, will also allow us to increase the matching point with our high-energy Regge description. The latter is a semi-local approach, i.e., it describes the amplitude on the average around a given energy [83] and is only to be used inside integrals. Indeed, in [54] a Regge fit to total cross sections in different $\pi\pi$ channels was used above 1.42 GeV, but no effort was made to match it to the partial-wave series within uncertainties. The reason was that in [54] the high-energy part only contributed as input for the integrals, and was relatively suppressed when testing or constraining the low-energy region of interest. However, since we now want a precise description also up to higher energies, better matching is needed to make sure there are no spurious artifacts [^2] We thank C. Hanhart for expressing interest and calling our attention to this issue.. Then, to improve the precision and consistency of our parameterizations, we will impose the FDR constraints, but now with a better matching and up to 1.61.61.6\,1.6GeV for the two FDRs containing P-wave contributions. This is almost 200200200\,200MeV higher than in [54, 68]. All in all, the whole global S0 wave or the other waves below 0.9 GeV will barely change from the CFD and global parameterizations provided in [54, 68], but there will be a clear improvement for the other waves at higher energies.
@@
Let us now describe the parameterizations used for each partial wave. We we will present our parameterizations and illustrate the dispersive constraints in terms of Solution I, which is the most popular and, as we will discover later on, seems to be slightly favored by the dispersive checks.
III Analytic parameterizations
@@
Customarily, the $\pi\pi\to\pi\pi$ partial-wave $S^{(I)}_{\ell}$-matrix element of definite isospin $I=0,1,2$ and angular momentum $\ell$, is parameterized in terms of two real functions:
@@
$t^{(I)}_{\ell}(s)=\frac{\eta^{(I)}_{\ell}(s)e^{2i\delta^{(I)}_{\ell}(s)}-1}{2i\sigma(s)},$ (3)
while the total amplitude $F^{(I)}(s,t)$ of definite isospin $I$ is normalized as follows:
@@
We will use several types of parameterizations for $t^{(I)}_{\ell}$, $\delta^{(I)}_{\ell}$, $\eta^{(I)}_{\ell}$, or $\cot\delta^{(I)}_{\ell}$. Some parameterizations will have physical features built in, such as factors to describe poles, zeros, peaks, or a specific threshold behavior, described in detail for each wave below. In this way, we try to mimic the appearance of some new effect at a given energy, like a new channel opening up with small uncertainties that grow at higher energies. However, other parameterizations will be purely phenomenological, in terms of the powers of some energy variables. They will be used in regions where we want to describe a relatively uniform error band and provide somewhat smaller correlations between their parameters. We will often use Chebyshev polynomials of the first kind, $p_{k}(x)$. They satisfy: $p_{0}(x)=1$, $p_{1}(x)=x$ and the rest are defined recursively in the interval $x\in[-1,1]$ as:
@@
Next, let us describe our global parameterizations, starting with the attractive P, D0, F, and G0 waves, all of which present some resonant behavior, followed by the repulsive S2, D2, and G2 waves. The indirect and minor changes of the S0 wave with respect to [68], are discussed in Appendix B.
III.1 P-wave parameterization
A global parameterization for this wave was already provided in [68]. Here we will improve it to allow its inelasticity to start at the $\pi\omega$ threshold at $\sqrt{s_{\pi\omega}}\equiv m_{\pi}+m_{\omega}=922.23\,$MeV.
In addition, this is the only wave where we will update the phase shift data in the elastic region below the $\pi\omega$ threshold. As explained before, this is because the dispersive analyses in [54] were not based on scattering data, but on a 2001 fit to the pion vector form factor [76]. Since then, many other data on the form factor have appeared, and we will therefore use as input the output of the pion vector form factor dispersive analysis by Colangelo et al. in [77], which we show as data points with uncertainties in the top panel of Fig. 1333 We thank C. Colangelo, M. Hoferichter and P. Stoffer for kindly providing us with their results.. Note that this analysis does not include the recent CMD-3 $e^{+}e^{-}\to\pi^{+}\pi^{-}$ data [84], which show sizable tension with previous results. Nevertheless, the impact of these data on the $\pi\pi$ P-wave phase shift analysis of [77] below the $\pi\omega$ threshold has been found to be relatively small [85], and therefore do not affect our results.
As stated above, we aim to provide simple formulae that can be easily implemented later for other purposes, while precisely describing the central value and uncertainty band. This last requirement makes it very difficult to provide a single parameterization covering the whole energy range when a wave has a lot of structure, as seen in the middle and bottom panels of Fig. 1. The data there come from [20], which we call Solution I here.
Therefore, our parameterization still has two pieces. The first one below 1.4 GeV, where the inelasticity is small ($\eta^{(1)}_{1}\geq 0.85$) and will be constrained by all FDRs (and Roy-like equations below 1.1 GeV). Note that this piece will now be a single analytic parameterization, whereas in [54] it also consisted of two pieces, above and below the inelastic threshold. The second part extends from 1.4 GeV to $\sim$1.8 GeV, where the inelasticity can be rather large at some energies, i.e. $\eta\sim 0.5$, and only two FDRs provide constraints up to 1.6 GeV. Above this energy, it is just an unconstrained fit to data. The applicability of the FDRs is dictated by the energy at which the partial-wave and Regge representations match within uncertainties.
Let us provide the detailed expressions of this parameterization.

### v2

# Global parametrizations of $\pi\pi$ scattering with dispersive constraints: Beyond the S0 wave

@@ Abstract
We provide new global parametrizations of $\pi\pi\to\pi\pi$ scattering for the S2, P, D, F, and G partial waves up to at least 1.8 GeV, easy to implement for phenomenological use. With earlier S0-wave parametrizations, slightly updated here, they reproduce previous partial wave dispersion analyses up to the $\pi\omega$ threshold. In addition, these new parametrizations have improved their description of recent P-wave data, the inelasticity in various waves, and their fulfillment of Roy-like and forward dispersion relations. The latter now test very high partial waves and have an improved matching with the Regge regime, extending those with P-wave contributions up to 1.6 GeV. Above 1.6 GeV and up to 1.8 GeV, or sometimes somewhat beyond, the parametrizations are simple unconstrained fits to data.
@@ I Introduction
Still, modern Hadron Physics demands precise and model-independent meson-meson scattering parametrizations. This requires the use of dispersion relations, which over roughly the last two decades have been successfully applied to describe scattering data on $\pi\pi$ [50, 51, 52, 53, 54, 55, 56, 57, 58, 59], but also on $\pi N$ [60, 61], $\gamma\gamma\to\pi\pi$ [62, 63], $\pi K$ [64, 65] and $\pi\pi\to K\bar{K}$ [66] (see [67] for a review of the last two).
@@
Unfortunately, such rigorous dispersive results, are not always suitable for later practical use. There are several reasons for this. First, they are often obtained numerically from integral equations, which makes them inconvenient for further phenomenological or experimental analyses. Second, they are frequently parametrized in the real axis by piecewise functions. They provide flexibility, but cannot be directly continued to the complex plane in search of poles. Finally, of the usual constraints used in the literature, partial-wave dispersion relations are in practice limited to energies around 111\,1GeV for $\pi\pi$ and $\pi K$ scattering, and to 1.41.41.4\,1.4 GeV for forward dispersion relations. Above these energies, partial-wave experimental data have not been tested nor described with dispersive constraints. In this work, we will partially alleviate these caveats.
@@
Recently, two of us with another collaborator [68] provided a set of relatively simple and ready-to-use “global” parametrizations that describe the scalar-isoscalar (S0) and vector-isovector (P1 or just P) partial waves of $\pi\pi\to\pi\pi$ scattering data up to somewhat below 2 GeV. They mimic the central values and uncertainties of the dispersive data analysis in [54], which was dispersively constrained up to 1.42 GeV. Beyond this energy, these parametrizations were three purely phenomenological fits to three different data sets from the literature. In addition, being analytic expressions, these parametrizations could be evaluated in the complex plane and they were made to reproduce the dispersive results, including the poles associated with the $\sigma/f_{0}(500)$, $f_{0}(980)$ and $\rho(770)$ resonances. Later, it was found in [69] that the poles of the $f_{0}(1370)$ and $f_{0}(1500)$ are also present in the dispersive analysis and global parametrization of the S0 wave. Furthermore, these global parametrizations are also consistent with the threshold parameters and the value of the S0-wave Adler zero in [54]. They have been widely used by both the theoretical and experimental hadron and particle physics communities.
@@
The aim of this work is threefold. First, extend this “global” parametrization approach beyond the partial waves S0 and P up to those with angular momentum $\ell=4$, trying to avoid piecewise parametrizations whenever possible. Second, improve the treatment of the P-wave inelasticity, allowing it to start at the $\pi\omega$ threshold. This was neglected in [54, 68] [^1] We thank G. Colangelo, M. Hoferichter, and P. Stoffer for pointing this to us., which focused more on the S0 wave and its resonances, but there is a considerable interest in the precise description of this wave. Third, extend the forward dispersion relation constraints to higher energies, while improving the matching with the high-energy Regge description. These improvements will have a relatively small effect on the S0 wave, whose global parameters will be slightly updated, but only because of the indirect effects of the changes in the other waves.
@@
The plan of this work is as follows: Sec. II presents the data to be described. In Sec. III we present the global partial-wave parametrizations. In Sec. IV we will revisit the dispersive constraints and obtain constrained Global Fits. We will discuss and summarize our results in Sec. V.
@@ II The input to be described
There are several data sets on $\pi\pi\to\pi\pi$ scattering partial waves $t^{(I)}_{\ell}(s)$, which almost reach a center-of-mass energy of $\sqrt{s}\sim 2\,$ GeV [20, 21, 22, 23, 24, 25, 26, 27, 28]. Here $I=0,1,2$ stands for isospin and $\ell=0,1,2,3,4...$ for angular momentum, although we will follow the usual spectroscopic notation and refer to them as S, P, D, F, G… waves, followed by their isospin number.
Unfortunately, these are indirect measurements, extracted as a sub-process of $\pi N\to\pi\pi N^{\prime}$ scattering, which is the observed reaction. These analyses assume the dominance of an almost on-shell one-pion exchange, along with other approximations. As a consequence of this indirect extraction, $\pi\pi$ data sets have large systematic uncertainties, and are often incompatible with each other, even when extracted from the same $\pi N\to\pi\pi N^{\prime}$ experiment. Furthermore, simple fits to individual data sets or to averaged sets do not comply well with dispersion relations [70, 71, 72, 73, 54]. This is why such relations were used to eliminate inconsistent data sets or as constraints to obtain a Constrained Fit to Data (CFD) [54]. In addition, these CFDs fulfill the normality requirements of the residual distribution [74], thus allowing for standard error-propagation methods. Consequently, the CFD parametrizations will be part of our input.
Why, then, not to use the CFD directly? Actually, it is perfectly fine to use them, although they were constructed with piecewise functions and only up to energies around 1.41.41.4\,1.4GeV. Moreover, the focus of [54] was the low-energy parameters and the region of the $\sigma/f_{0}(500)$, $f_{0}(980)$, and $\rho(770)$ resonances. The description above 1.1 GeV, or of higher waves, was provided primarily to serve as input for the integral representation of the S0 and P amplitudes below that energy. The need for “global” parametrizations is justified in that many applications require a wider energy range and waves with higher angular momentum, and because the CFD piecewise expressions cannot be straightforwardly continued to the complex plane to describe resonance poles. This is why in [68] some of us provided such global parametrizations for the S0 and P waves. Here we extend this approach to six more waves: S2, D0, D2, F, G0, and G2. In addition, we improve some details of the P wave above 0.9 GeV, and expand the dispersive constraints further than 1.4 GeV for all the waves. Moreover, above roughly 1 GeV, we improve the precision of all waves beyond the S0.
@@
Hence, we will use as input the threshold parameters obtained from the CFD parametrizations in [54] or sum rules [73]. In addition, we will consider the CFD in the real $s$ axis as input for the S2 and D waves, although only below 0.9 GeV, as we plan to improve the treatment of their inelasticities, which in [54] were only taken into account above the $K\bar{K}$ threshold.
@@
Concerning the S0 wave for energies above 1.4 GeV, “Solution b” and “Solution (- - -)” are very incompatible. Other solutions were already disfavored in the very same CERN-Munich 1975 analysis, although the “(- + -) solution” has recently been resurrected [78]. Nevertheless, all these solutions have caveats (see [79, 70, 45] for detailed discussions). In particular, at high energies, the elastic cross section of “Solution b” is larger than the inelastic one, which is at odds with the observations in $\pi N$, $KN$, and $NN$ scattering, and there is no obvious reason why $\pi\pi$ scattering should be different. In addition, right above 1 GeV, the S0-wave data of the other CERN-Munich solutions prefer less inelasticity, i.e. the so-called “non-dip” scenario [54], which is disfavored by Roy-like dispersive constraints [54, 56]. Later on, we will remove such data points for these solutions. Moreover, if the inelasticity is large, the solution in terms of phase shift and elasticity is not unique [80, 81]. “Solution b” is an example of an almost elastic case and “Solution (- - -)” is strongly inelastic. This, of course, applies to other waves. Furthermore, a slight modification of the “(- + -) solution” by one of the members of the CERN-Munich collaboration [78] considering an inelastic S2 wave, seems consistent with the dispersive representation [68], finding some qualitative agreement with the GAMS experiment on $\pi^{-}p\rightarrow\pi^{0}\pi^{0}n$ [82]. Such a “(- + -) solution” differs from the “b” and “(- - -)” solutions only above 1.4 GeV and hints at the presence of the $f_{0}(1500)$ resonance, not evident in other solutions. Last but not least, the convergence of the partial-wave expansion is questionable, since already at 1.7 GeV the F wave is as large as the P wave, the D0 as the S0, and the D2 is larger than the S2.
@@
Not only the S0 but also the P, D0, and F waves, have three solutions, although they are not so different among themselves. The existence of different data sets leads to three different constrained global fits in [68], almost identical up to 1.43 GeV. Namely, above 1.43 GeV: i) “Solution I” describes the data of [20, 25, 28], ii) “Solution II” describes the (- - -) data of [26], and iii) “Solution III” describes the data of the “(- + -) solution” updated in [78]. We will follow a similar strategy here, considering the three sets of data called Solutions I, II, and III again, and their corresponding fits that we will call Global Fits I, II, and III. There are no scattering data for the G0 wave and we will build an educated guess from other information.
@@
Concerning the $I=2$ waves, we also use CFD input below 0.9 GeV for the repulsive S2 and D2, often called “exotic waves”. No CFD was available for the G2 wave, which is also exotic. Above that energy we consider the 1973 data from the Rochester collaboration measured at Brookhaven (Cohen et al. [23]), the Paris-Bari collaboration obtained in 1973 at CERN (Durusoy et al. [21]), the 1974 CERN-Saclay data (Losty et al. [22]) and the 1977 Amsterdam-CERN-Munich data (Hoogland et al. [27]). These experiments are roughly in agreement, partly because these waves are small and the uncertainties become comparatively large. Thus, we do not contemplate alternative data sets for these waves and there will be only one unconstrained fit for each $I=2$ wave. However, since the dispersive constraints affect all waves simultaneously, the alternative data sets from other waves will indirectly affect the $I=2$ waves, which will also have three different constrained fits, although they will look rather similar in practice.
@@
Having fits for all these waves up to at least 1.8 GeV, and sometimes even up to 2.1 GeV, will also allow us to increase the matching point with our high-energy Regge description. The latter is a semi-local approach, i.e., it describes the amplitude on the average around a given energy [83] and is only to be used inside integrals. Indeed, in [54] a Regge fit to total cross sections in different $\pi\pi$ channels was used above 1.42 GeV, but no effort was made to match it to the partial-wave series within uncertainties. The reason was that in [54] the high-energy part only contributed as input for the integrals, and was relatively suppressed when testing or constraining the low-energy region of interest. However, since we now want a precise description also up to higher energies, better matching is needed to make sure there are no spurious artifacts [^2] We thank C. Hanhart for expressing interest and calling our attention to this issue.. Then, to improve the precision and consistency of our parametrizations, we will impose the FDR constraints, but now with a better matching and up to 1.61.61.6\,1.6GeV for the two FDRs containing P-wave contributions. This is almost 200200200\,200MeV higher than in [54, 68]. All in all, the whole global S0 wave or the other waves below 0.9 GeV will barely change from the CFD and global parametrizations provided in [54, 68], but there will be a clear improvement for the other waves at higher energies.
@@
Let us now describe the parametrizations used for each partial wave. We will present our parametrizations and illustrate the dispersive constraints in terms of Solution I, which is the most popular and, as we will discover later on, seems to be slightly favored by the dispersive checks.
III Analytic parametrizations
@@
Customarily, the $\pi\pi\to\pi\pi$ partial-wave $S^{(I)}_{\ell}$-matrix element of definite isospin $I=0,1,2$ and angular momentum $\ell$, is parametrized in terms of two real functions:
@@
$t^{(I)}_{\ell}(s)=\frac{\eta^{(I)}_{\ell}(s)e^{2i\delta^{(I)}_{\ell}(s)}-1}{2i\sigma(s)}.$ (3)
Note that we are not describing the full $S$-matrix, but just its first diagonal element $S_{11}$, where $1=\pi\pi$ (we have suppressed the spin and isospin indices momentarily). On this single element, unitarity above the $\pi\pi$ threshold only implies that $\eta(s)=|S_{11}|\leq 1$.
In contrast, in the popular coupled channel approach, one describes all the $S_{ij}$-matrix elements. For that, one has to choose a priori how many channels $i,j=1,...n$ to consider and to identify each of them (and their thresholds). Within that approach the elasticity $\eta(s)\equiv|S_{11}|=\sqrt{1-|{S_{12}}|^{2}....-|{S_{1n}}|^{2}}$, and thus it is possible to know the partial contribution of each coupled channel to the inelasticity of $S_{11}$. This is impossible in our approach since we only determine the total elasticity. On the positive side, our approach avoids the model dependence due to deciding what channels are open or not and how to describe them, which, for multiple-meson states, is usually done in strongly model-dependent quasi-two-body approximations.
In practice, to fit the data, we have to decide where we allow the inelasticity to set in. In certain channels, the choice is easy because the inelasticity data appears right after a prominent two-body threshold like $K\bar{K}$, $\omega\pi$. However, multibody thresholds like $4\pi$ or $\pi\pi K\bar{K}$ or six pions, etc., open up more smoothly (but in coupled channel approaches are frequently modeled as $\rho\rho$, $\sigma\sigma$, $\rho\pi\pi$, …). When inelasticity is observed with no obvious two-body threshold right below, we have allowed it to open up at some effective threshold below the inelastic data. But remember that once we have a non-vanishing total inelasticity, it contains all possible inelastic contributions.
This said, we return to our previous notation and define the total amplitude $F^{(I)}(s,t)$ of definite isospin $I$, normalized as follows:
@@
We will use several types of parametrizations for $t^{(I)}_{\ell}$, $\delta^{(I)}_{\ell}$, $\eta^{(I)}_{\ell}$, or $\cot\delta^{(I)}_{\ell}$. Some parametrizations will have physical features built in, such as factors to describe poles, zeros, peaks, or a specific threshold behavior, described in detail for each wave below. In this way, we try to mimic the appearance of some new effect at a given energy, like a new channel opening up with small uncertainties that grow at higher energies. However, other parametrizations will be purely phenomenological, in terms of the powers of some energy variables. They will be used in regions where we want to describe a relatively uniform error band and provide somewhat smaller correlations between their parameters. We will often use Chebyshev polynomials of the first kind, $p_{k}(x)$. They satisfy: $p_{0}(x)=1$, $p_{1}(x)=x$ and the rest are defined recursively in the interval $x\in[-1,1]$ as:
@@
Next, let us describe our global parametrizations, starting with the attractive P, D0, F, and G0 waves, all of which present some resonant behavior, followed by the repulsive S2, D2, and G2 waves. The indirect and minor changes of the S0 wave with respect to [68], are discussed in Appendix B.
III.1 P-wave parametrization
A global parametrization for this wave was already provided in [68]. Here we will improve it to allow its inelasticity to start at the $\pi\omega$ threshold at $\sqrt{s_{\pi\omega}}\equiv m_{\pi}+m_{\omega}=922.23\,$MeV.
In addition, this is the only wave where we will update the phase shift data in the elastic region below the $\pi\omega$ threshold. As explained before, this is because the dispersive analyses in [54] were not based on scattering data, but on a 2001 fit to the pion vector form factor [76]. Since then, many other data on the form factor have appeared, and we will therefore use as input the output of the pion vector form factor dispersive analysis by Colangelo et al. in [77], which we show as data points with uncertainties in the top panel of Fig. 1.333 We thank C. Colangelo, M. Hoferichter and P. Stoffer for kindly providing us with their results. Note that this analysis does not include the recent CMD-3 $e^{+}e^{-}\to\pi^{+}\pi^{-}$ data [84], which show sizable tension with previous results. Nevertheless, the impact of these data on the $\pi\pi$ P-wave phase shift analysis of [77] below the $\pi\omega$ threshold has been found to be relatively small [85, 86], and therefore do not affect our results.
As stated above, we aim to provide simple formulae that can be easily implemented later for other purposes, while precisely describing the central value and uncertainty band. This last requirement makes it very difficult to provide a single parametrization covering the whole energy range when a wave has a lot of structure, as seen in the middle and bottom panels of Fig. 1. The data there come from [20], which we call Solution I here.
Therefore, our parametrization still has two pieces. The first one below 1.4 GeV, where the inelasticity is small ($\eta^{(1)}_{1}\geq 0.85$) and will be constrained by all FDRs (and Roy-like equations below 1.1 GeV). Note that this piece will now be a single analytic parametrization, whereas in [54] it also consisted of two pieces, above and below the inelastic threshold. The second part extends from 1.4 GeV to $\sim$1.8 GeV, where the inelasticity can be rather large at some energies, i.e. $\eta\sim 0.5$, and only two FDRs provide constraints up to 1.6 GeV. Above this energy, it is just an unconstrained fit to data. The applicability of the FDRs is dictated by the energy at which the partial-wave and Regge representations match within uncertainties.
Let us provide the detailed expressions of this parametrization.